OpenStudy (anonymous):

What is the value of the system determinant for the following system of equations? 6x + y = 4 5x + y - 7 = 0 -1 1 31

OpenStudy (anonymous):

find the matrix first

OpenStudy (anonymous):

6 1 -4 5 1 -7 = 6[(-7) -(-4)]-[-42-(-20)] +(-4)[6-5] =6(-3)-(-22)-4(1) =-18+22-4 =0

OpenStudy (anonymous):

ur determinant is 0